OpenStudy (anonymous):

Evaluate the function rule for the given value. f(x) = 3^x for x = –5 a.1/243 b. –125 c. 1/729 d. –15

OpenStudy (anonymous):

Put x = -5 in function..

OpenStudy (anonymous):

\[f(-5) = 3^{-5} =??\]

OpenStudy (anonymous):

And use: \[a^{-b} = \frac{1}{a^b}\] a = 3 and b = 5 here..

OpenStudy (anonymous):

given f(x)=3^x so,f(-5)=3^-5 =1/3^5 =1/243

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome =)